A chemist needs to store a liquid inside a container in a laboratory room. She is trying to determine whether to store the liquid in a regular square pyramid or in a right circular cone storage container. The pyramid container has a base that is 6 inches by 6 inches and a height of 8 inches. The cone container has a base with a diameter of 6 inches and a height of 10 inches. Which container --the pyramid or the cone -- has a greater volume, and by how many cubic inches?
Answer by mananth(16946)

Volume of square pyramid = a^2 * h/3
a= 6*6=36 in^2
h=8 in
V= 36*8/3 = 96 in^3
Volume of cone = 1/3 pi r^2 h
r=3, h= 10 in
V= 1/3 * pi * 3^2 *10 =94.25 in^3
The square pyramid container has a greater volume by 96−94.25=1.75 in^3